in a recent survey of 25 voters, 17 favor a new city regulation and 8 oppose it. what is the probability that in a random sample of 6 respondents from this survey, exactly 2 favor the proposed regulation 4 oppose it?
Answer by stanbon(75887)

in a recent survey of 25 voters, 17 favor a new city regulation and 8 oppose it. what is the probability that in a random sample of 6 respondents from this survey, exactly 2 favor the proposed regulation 4 oppose it?
------------------
Binomial Problem with n=6; p = 17/25 ; x = 2
---
P(x= 2) = 6C2*(17/25)^2(8/25)^4 = binompdf(6,17/25,2) = 0.0727
=====================
I used a TI calculator to get the result.
=====================
Cheers,
Stan H.